# Surrogate models for lightcurves and photosphere properties of Type II supernovae

### v1

@@ Abstract
Inferences on the properties Type II supernovae (SNe) can provide significant insights into the lives and deaths of the astrophysical population of massive stars and potentially provide measurements of luminosity distance, independent of the distance ladder. Here, we introduce surrogate models for the photospheric properties and lightcurves of Type II SNe trained on a large grid of simulations from the radiation hydrodynamics code, stella. The trained model can accurately and efficiently ($\sim 30$ms) predict the lightcurves and properties of Type II SNe within a large parameter space of progenitor ($10-18M_{\odot}$ at ZAMS) and nickel masses ($0.001-0.3M_{\odot}$), progenitor mass-loss rate ($10^{-5}-10^{-1}~{}M_{\odot}$yr-1), CSM radius ($1-10\times 10^{14}$cm), and SN explosion energies ($0.5-5\times 10^{51}$erg). We validate this model through inference on lightcurves and photosphere properties drawn directly from the original stella simulations not included in training. In particular, for a synthetic Type II SNe observed within the 10-year LSST survey, we find we can measure the progenitor and nickel masses with $\approx 9\%$ and $\approx 25\%$ precision, respectively, when fitting the photometric data while accounting for the uncertainty in the surrogate model itself. Meanwhile, from real observations of SN 2004et, SN 2012aw, and SN 2017gmr we infer a progenitor ZAMS mass of $12.15_{-1.06}^{+1.03}M_{\odot}$, $10.61_{-0.32}^{+0.37}M_{\odot}$, $10.4\pm 0.3M_{\odot}$, respectively. We discuss systematic uncertainties from our surrogate modelling approach and likelihood approaches to account for these uncertainties. We further discuss future extensions to the model to enable stronger constraints on properties of Type II SNe and their progenitors, for cosmological applications, and applications of our surrogate modelling approach to other transients.
@@ 1 Introduction
Lightcurve and spectral modelling of past Type II SNe has yielded insights into the explosion energy, nucleosynthetic yields and progenitor properties of massive stars (Förster et al., 2018; Davies & Beasor, 2018; Goldberg et al., 2019; Martinez et al., 2022; Subrayan et al., 2023; Silva-Farfán et al., 2024; Hinds et al., 2025). These properties have been used to elucidate the explosion mechanism (e.g., Burrows & Vartanyan, 2021) and facilitated the use of Type IIP SNe as a new distance indicator for cosmology (Hamuy & Pinto, 2002; de Jaeger et al., 2020). As the dataset continues to grow through time-domain surveys, we are increasingly in need of computationally efficient models to extract physical parameters from often sparsely sampled photometric observations. While detailed numerical simulations provide the most physically grounded models (e.g., Bersten et al., 2011; Dessart et al., 2013; Moriya et al., 2023), their significant computational expense prohibits direct application for large-scale inference for hundreds or thousands of events. Meanwhile, simpler analytical or semi-analytical models (Popov, 1993; Nagy & Vinkó, 2016), while faster, often rely on simplifying assumptions that limit their ability to adequately capture the complexity revealed by detailed simulations. This is especially true in cases where the lightcurve may be powered by CSM interaction or behave differently due to different assumptions on 56Ni mixing (e.g., Förster et al., 2018; Bruch et al., 2021).
Moriya et al. (2023) recently published a large grid of synthetic Type II SN models based on red supergiant progenitors (RSG) from Sukhbold et al. (2016), including properties such as the photospheric temperatures and radius, bolometric luminosities, as well as spectrum computed through the radiation hydrodynamics code  STELLA (Blinnikov et al., 1998, 2000, 2006). This grid, encompassing 228,016 models, systematically explores a wide parameter space, including progenitor zero age main sequence (ZAMS) mass, explosion energy, 56Ni mass, and various wind-like CSM properties (mass-loss rate, extent, structure). Comparisons with this grid have already facilitated a wide variety of analyses into the properties of Type II SNe (e.g., Moriya & Singh, 2024; Chen et al., 2025; Hinds et al., 2025). However, while this grid forms a crucial basis for systematic studies, direct interpolation within such a high-dimensional grid for parameter inference can still be challenging, preventing the use of these models to directly fit lightcurves and extracting posteriors on explosion and progenitor properties.
@@
Here, we introduce a new surrogate models for the spectra, lightcurves and photosphere properties from this grid, that can be used to efficiently and accurately interpolate the outputs from the numerical simulations. In particular, in Sec. 2, we describe the design, training process, and performance of our surrogate models. In Sec. 3, we validate our surrogate models by performing inference on bolometric luminosity lightcurve and photosphere properties against the original stella simulations, as well as for synthetic photometric observations with the Vera Rubin Observatory. We then demonstrate the utility of the model by inferring the physical parameters of a few Type II SNe by fitting their photometric observations in Sec. 4. Finally, we discuss future extensions to our models, systematic uncertainties associated with the surrogate modelling approach and conclude in Sec. 5.
@@ 2 Surrogate model
2.1 Training data and model design
As mentioned above, Moriya et al. (2023) provide us a total of $228,015$ different simulations [^1] We note that one simulation from the original 228,016 simulations is missing bolometric luminosity data and could not be run again.. Each simulation contains data for the bolometric luminosity, photosphere temperature, photosphere radius, and spectral energy distribution (SED) at $100$ wavelengths from $1-50000$ at multiple epochs in time. Ideally, a single surrogate model trained directly on the SED can be used to reconstruct photometry and bolometric luminosities. Moreover, it can capture the effects of any emission or absorption lines in the spectrum, enabling fitting directly to spectroscopic data. In practice, the SEDs in our training data do not include any spectral features, they are also not sufficiently densely sampled to enable detailed inferences against high-resolution spectra. Therefore, we build multiple independent surrogates for the bolometric luminosity, photosphere properties (temperature and radius), and for the spectrum itself which can be used to generate photometry. This approach provides flexibility enabling us to perform inference on different types of data (or combine multiple diagnostics) which can be used to further test models. We note that the photosphere temperature and radius could also be used to generate photometry assuming a blackbody SED, but Type II SNe are not well approximated by a blackbody at late phases or at ultra-violet (UV) wavelengths, so we do not recommend this approach.
@@ 2.1.1 Luminosity and photosphere property surrogates
Our first step towards building our surrogate models was to prepare the training data, as each simulation output is on different time samples, we processed all the different outputs (bolometric luminosity, photosphere temperature, and photosphere radius) for all simulations onto a consistent time array from 0.1 to 400 days (in rest frame time). We do this through linear interpolation as the simulation outputs are densely sampled and smooth functions. We then randomly split the $228,015$ simulations into a testing and training set of $30000$ and $198015$ simulations, with the latter further split into a $80-20\%$ training and testing set for training and inter-epoch validation. We also transform the input parameters (summarised in Table 1) and outputs onto a unit Gaussian to aid training.
@@
$M_{\rm ZAMS}$ [$M_{\odot}$] Initial ZAMS mass of the progenitor 10 - 18
56Ni [$M_{\odot}$] Mass of nickel-56 0.001-0.3
$\dot{M}$ [$M_{\odot}$yr-1] Mass loss rate of the progenitor $10^{-5}-10^{-1}$
@@
$R_{\rm csm}$ [$10^{14}$cm] Radius of the circumstellar material 1-10
$E_{\rm sn}$ [$10^{51}$erg] Energy of the supernova explosion 0.5-5
@@
As all simulations are all on discrete points, e.g., $10,12,14,16,18\ M_{\odot}$, we also augment our training data with additional samples that are small scatters around the simulation input and outputs. We find this additional scatter enables the model to generalise far more reliably. We then employ a feed forward neural network via tensorflow (Abadi et al., 2016) with a combination of up to 5 layers with a hyperbolic tangent activation function and a combination of mean absolute and mean squared error loss functions. We train with a batch size of 5000 (to fully leverage GPUs) and for up to 400 epochs (with early stopping, in the case the model does not improve for more than 10 epochs). Our total training time is $\approx 10$ mins on a M1 pro Macbook GPU, with the output being three different trained models which independently predict the bolometric luminosity and photosphere temperature and radius.
@@ 2.1.2 Spectrum surrogate
To create a surrogate model for the spectrum we use a combination of an autoencoder, a principle-component analysis, and a feed-forward neural network. We first preprocess our spectrum data onto uniform arrays in time (geometrically from $0.1-400$ days) and wavelength(geometrically from $500-49500$ using a cubic interpolator. We have verified that this faithfully captures the true spectrum across a range of simulations. We also normalise the flux data such that all values are between $0-1$ and augment our data as described above. The autoencoder encodes our uniform high-dimensional flux data into a $64$ dimensional latent-space, while also training a decoder to invert this process. We further reduce the dimensionality by performing a principal component analysis to $32$ dimensions with minimal impact on our accuracy. We then train a feed-forward neural network (with similar design to our surrogates above) to learn the mappings from the $6$ physical input parameters onto the reduced $32$ dimensional latent space. The full pipeline takes $\approx 40$ mins to train with adapting batch sizes (to ensure more granular gradient updates towards the end of training) and a total of 100 epochs. Our surrogate model can then be used to generate the spectrum for any arbitrary combination of parameters (within the ranges specified in Table 1), we can then generate photometry by integrating the spectrum over the bandpass of any filter via Redback (Sarin et al., 2024) utilising filter transmission curves through the Spanish Virtual Observatory (Rodrigo & Solano, 2020).
@@ 2.2.1 Luminosity and photosphere property surrogates
To evaluate the performance of our trained models, we consider a range of metrics relative to an unseen training data of $30000$ simulations. Our first consideration is accuracy, across the training data, we see mean squared errors of $0.0056$, $0.06$, and $0.01$ for the luminosity, temperature, and radius surrogates, respectively, on the normalized values. Moreover, our models are also extremely fast, with each prediction taking $\approx 30$ms. Furthermore, due to the design of a neural network, we can make predictions for large batches of simulations at once, for example, all outputs for the $30000$ test simulations are generated within $\approx 30$ms, which can provide significant benefits for inference studies as we will discuss later. In Fig. 1, we plot a random set of bolometric luminosities, photosphere radius and temperature from our training data (solid curves) alongside the prediction of the surrogate model (dashed curves). The bottom panel indicates the normalised residual, which highlights that across the bulk of times, our normalized residuals are smaller than $5\%$, this is smaller than the numerical noise in the simulations itself, particularly at early times.
We further investigate the residuals across our test data to see if there are any patterns of failure with our surrogate model. In particular, in Fig. 2, we show $1\sigma$ distribution on the normalised residuals as a function of time from the full sample of our $30000$ testing data. First, we draw attention to the luminosity residuals, which indicates a discrepancy between our surrogate model and the testing data of as large as $15\%$ at $t\sim 1$d, which could influence our results when performing inference if unaccounted for. Our residuals also highlight an unusual peak at $\approx 100$days for all models, this is likely a product of a numerical noise in some simulations at these epochs. However, as this is only on the $\approx 5\%$ level, we consider this accurate for our purposes of inferences, as opposed to the larger error in the bolometric luminosity surrogate at early times ($t\lesssim 1$day). We note that while above we are showing the $1\sigma$ credible interval, our median normalised residuals are significantly smaller and on the order of $10^{-3}$, highlighting strong consistency between the surrogate model predictions and the training data.
Figure 3: $90\%$ credible interval (shown by blue shaded region) of the bolometric luminosity, photosphere temperature and radius generated from $50000$ random samples our prior. We also show the predicted properties for a Type II SN with $M_{\rm ZAMS}=13M_{\odot}$ (red curves), not included in our training or simulated data.

### v2

@@ Abstract
Inferences on the properties Type II supernovae (SNe) can provide significant insights into the lives and deaths of the astrophysical population of massive stars and potentially provide measurements of luminosity distance, independent of the distance ladder. Here, we introduce surrogate models for the photospheric properties and lightcurves of Type II SNe trained on a large grid of simulations from the radiation hydrodynamics code, stella. The trained model can accurately and efficiently ($\sim 30$ms) predict the lightcurves and properties of Type II SNe within a large parameter space of progenitor ($10-18M_{\odot}$ at ZAMS) and nickel masses ($0.001-0.3M_{\odot}$), progenitor mass-loss rate ($10^{-5}-10^{-1}~M_{\odot}$yr-1), CSM radius ($1-10\times 10^{14}$cm), and SN explosion energies ($0.5-5\times 10^{51}$erg). We validate this model through inference on lightcurves and photosphere properties drawn directly from the original stella simulations not included in training. In particular, for a synthetic Type II SNe observed within the 10-year LSST survey, we find we can measure the progenitor and nickel masses with $\approx 9\%$ and $\approx 25\%$ precision, respectively, when fitting the photometric data while accounting for the uncertainty in the surrogate model itself. Meanwhile, from real observations of SN 2004et, SN 2012aw, and SN 2017gmr we infer a progenitor ZAMS mass of $12.15_{-1.06}^{+1.03}M_{\odot}$, $10.61_{-0.32}^{+0.37}M_{\odot}$, $10.4\pm 0.3M_{\odot}$, respectively. We discuss systematic uncertainties from our surrogate modelling approach and likelihood approaches to account for these uncertainties. We further discuss future extensions to the model to enable stronger constraints on properties of Type II SNe and their progenitors, and applications of our surrogate modelling approach to other transients.
@@ 1 Introduction
Lightcurve and spectral modelling of past Type II SNe has yielded insights into the explosion energy, nucleosynthetic yields and progenitor properties of massive stars (Förster et al., 2018; Davies & Beasor, 2018; Goldberg et al., 2019; Martinez et al., 2022; Subrayan et al., 2023; Silva-Farfán et al., 2024; Hinds et al., 2025). These properties have been used to elucidate the explosion mechanism (e.g., Burrows & Vartanyan, 2021). As the dataset continues to grow through time-domain surveys, we are increasingly in need of computationally efficient models to extract physical parameters from often sparsely sampled photometric observations. While detailed numerical simulations provide the most physically grounded models (e.g., Bersten et al., 2011; Dessart et al., 2013; Moriya et al., 2023), their significant computational expense prohibits direct application for large-scale inference for hundreds or thousands of events. Meanwhile, simpler analytical or semi-analytical models (Popov, 1993; Nagy & Vinkó, 2016), while faster, often rely on simplifying assumptions, such as uniform density profiles, spherical symmetry, constant opacity, and simplified energy deposition from radioactive decay. These simplifications limit their ability to adequately capture the complexity revealed by detailed simulations and probed by detailed observations. This is especially true in cases where the lightcurve may be powered by CSM interaction or behave differently due to different assumptions on 56Ni mixing (e.g., Förster et al., 2018; Bruch et al., 2021).
Moriya et al. (2023) recently published a large grid of synthetic Type II SN models based on red supergiant progenitors (RSG) from Sukhbold et al. (2016), including properties such as the photospheric temperatures and radius, bolometric luminosities, as well as spectrum computed through the radiation hydrodynamics code stella (Blinnikov et al., 1998, 2000, 2006). The grid from Moriya et al. (2023), encompassing 228,016 models, systematically explores a wide parameter space, including progenitor zero age main sequence (ZAMS) mass, explosion energy, 56Ni mass, and various wind-like CSM properties (mass-loss rate, extent, structure). Comparisons with this grid have already facilitated a wide variety of analyses into the properties of Type II SNe (e.g., Moriya & Singh, 2024; Chen et al., 2025; Hinds et al., 2025). However, while this grid forms a crucial basis for systematic studies, direct interpolation within such a high-dimensional grid for parameter inference can still be challenging, preventing the use of these models to directly fit lightcurves and extracting posteriors on explosion and progenitor properties.
@@
Here, we introduce a new surrogate models for the spectra, lightcurves and photosphere properties from this grid, that can be used to efficiently and accurately interpolate the outputs from the numerical simulations. In particular, in Sec. 2, we describe the design, training process, and performance of our surrogate models. In Sec. 3, we validate our surrogate models by performing inference on bolometric luminosity lightcurve and photosphere properties against the original stella simulations, as well as for synthetic photometric observations with the Vera Rubin Observatory. We then demonstrate the utility of the model by inferring the physical parameters of a few Type II SNe by fitting their photometric observations in Sec. 4. Finally, we discuss future extensions to our models, systematic uncertainties associated with the surrogate modelling approach, and conclude in Sec. 5.
@@ 2 Surrogate model
2.1 Training grid assumptions
As mentioned above, Moriya et al. (2023) provide us a total of $228,015$ different simulations [^1] We note that one simulation from the original 228,016 simulations is missing bolometric luminosity data and could not be run again.. We briefly describe the primary ingredients, modelling assumptions, and limitations below, but refer the interested reader to Moriya et al. (2023); Moriya & Singh (2024) for further details. All parameters of the model, including their description and ranges are summarised in Table 1.
The model adopts red supergiant progenitors from Sukhbold et al. (2016) and attach a CSM of the form
$\rho_{\rm CSM}(r)=\frac{\dot{M}}{4\pi v_{\rm wind}(r)}r^{-2},$ (1)
where the velocity, $v\propto(1-R_{0}/r)^{\beta}$, where $R_{0}$ is the progenitor radius and $\beta$ sets the steepness of the density profile. This CSM structure extends out to $R_{\rm CSM}$, with the maximum value of $R_{\rm CSM}=10^{15}$cm motivated by observations (Moriya et al., 2023). The model then further assumes some explosion energy $E_{\rm SN}$ and a mass of 56Ni that is uniformly mixed up to the half mass of the hydrogen-rich envelope. Lightcurves and photosphere properties are then generated using stella (Blinnikov et al., 1998, 2000, 2006).
In general, this setup is well suited to capture the broadband properties of Type II SNe. However, there are a number of limitations with this setup which could affect our conclusions when applied to real data. As discussed in Moriya & Singh (2024), the choice of progenitor model and parameterisation has an impact on the constraints on the progenitor mass due to the strong correlation of this parameter with $R_{0}$, which is implicitly fixed by the choice of progenitor model. Moreover, as the radius out to which 56Ni is distributed is fixed across all simulations, we can not capture any discrepancies in the lightcurve due to this effect alone. Furthermore, at late-times as the supernova becomes nebular, the local-thermodynamic equilibrium (LTE) assumption built into stella breaks down, making it unsuitable for estimating the properties of Type II SNe at $\gtrsim 100$ days post explosion.
2.2 Training data and surrogate model design
Each simulation from Moriya et al. (2023) contains data for the bolometric luminosity, photosphere temperature, photosphere radius, and spectral energy distribution (SED) at $100$ wavelengths from 1500001-50000 Å at multiple epochs in time. Ideally, a single surrogate model trained directly on the SED can be used to reconstruct photometry and bolometric luminosities. Moreover, it can capture the effects of any emission or absorption lines in the spectrum, enabling fitting directly to spectroscopic data. In practice, the SEDs in our training data do not include any spectral features, they are also not sufficiently densely sampled to enable detailed inferences against high-resolution spectra. Therefore, we build multiple independent surrogates for the bolometric luminosity, photosphere properties (temperature and radius), and for the spectrum itself which can be used to generate photometry. This approach provides flexibility enabling us to perform inference on different types of data (or combine multiple diagnostics) which can be used to further test models. We note that the photosphere temperature and radius could also be used to generate photometry assuming a blackbody SED, but Type II SNe are not well approximated by a blackbody at late phases or at ultra-violet (UV) wavelengths, so we do not recommend this approach.
@@ 2.2.1 Luminosity and photosphere property surrogates
Our first step towards building our surrogate models was to prepare the training data, as each simulation output is on different time samples, we processed all the different outputs (bolometric luminosity, photosphere temperature, and photosphere radius) for all simulations onto a consistent time array from 0.1 to 400 days (in rest frame time). We do this through linear interpolation as the simulation outputs are densely sampled and smooth functions.
We randomly split the $228,015$ simulations into a testing set ($30,000$ simulations) and training set ($198,015$ simulations). The training set is further divided into training ($158,412$ simulations, i.e., $80\%$ of the training set) and validation ($39,603$ simulations, i.e., $20\%$ of the training set) subsets for model development and hyperparameter tuning. We also transform the input parameters and outputs onto a unit Gaussian to aid training. The validation simulations are used during each epoch of training to validate the model that is trained on the training set. The testing data is used after training is finished to test the final trained model.
@@
$_{\rm ZAMS} [MM_{\odot}$] Initial ZAMS mass of the progenitor 10 - 18
56Ni $M_{\odot}$] Mass of nickel-56 0.001-0.3
$\dot{M}$ $M_{\odot}$yr-1] Mass loss rate of the progenitor $10^{-5}-10^{-1}$
@@
$R_{\rm csm}$ $10^{14}$cm] Radius of the circumstellar material 1-10
$E_{\rm sn}$ $10^{51}$erg] Energy of the supernova explosion 0.5-5
@@
As all simulations are on discrete points, e.g., $10,12,14,16,18\ M_{\odot}$ for the ZAMS mass, we also augment our training data by adding Gaussian noise ($\sigma=0.01$ in normalized parameter space) to both inputs and outputs for each simulation, effectively creating 5 additional samples per original simulation. This augmentation increases our training set size to $\approx 800,000$ samples. We then employ a feed forward neural network via tensorflow (Abadi et al., 2016) with a combination of up to 5 layers with a hyperbolic tangent activation function and a combination of mean absolute and mean squared error loss functions. We train with a batch size of 5000 (to fully leverage GPUs) and for up to 400 epochs (with early stopping, in the case the model does not improve for more than 10 epochs). Our total training time is $\approx 10$ mins on a M1 pro Macbook GPU, with the output being three different trained models which independently predict the bolometric luminosity and photosphere temperature and radius. The augmentation has a significant impact on the accuracy of our trained model on the unseen testing set, improving loss by $\gtrsim 5\%$, while also enabling the model to generalize more reliably between discrete grid points.
@@ 2.2.2 Spectrum surrogate
To create a surrogate model for the spectrum we use a combination of an autoencoder, a principle-component analysis, and a feed-forward neural network. We first preprocess our spectrum data onto uniform arrays in time (geometrically from 0.14000.1-400 days) and wavelength(geometrically from 50049500500-49500 Å using a cubic interpolator. We have verified that this faithfully captures the true spectrum across a range of simulations. We also normalise the flux data such that all values are between 010-1 and augment our data as described above. The autoencoder encodes our uniform high-dimensional flux data into a $64$ dimensional latent-space, while also training a decoder to invert this process. We further reduce the dimensionality by performing a principal component analysis to $32$ dimensions with minimal impact on our accuracy. We then train a feed-forward neural network (with similar design to our surrogates above) to learn the mappings from the $6$ physical input parameters onto the reduced $32$ dimensional latent space. The full pipeline takes $\approx 40$ mins to train with adapting batch sizes (to ensure more granular gradient updates towards the end of training) and a total of 100 epochs. Our surrogate model can then be used to generate the spectrum for any arbitrary combination of parameters (within the ranges specified in Table 1), we can then generate photometry by integrating the spectrum over the bandpass of any filter via Redback (Sarin et al., 2024) utilising filter transmission curves through the Spanish Virtual Observatory (Rodrigo & Solano, 2020).
Figure 3: $90\%$ credible interval (shown by blue shaded region) of the bolometric luminosity, photosphere temperature and radius generated from $50000$ random samples our prior. We also show the predicted properties for a Type II SN with $M_{\rm ZAMS}=13~M_{\odot}$, $^{56}\rm{Ni}=0.02~M_{\odot}$, $\dot{M}=10^{-3.1}$ $M_{\odot}$yr-1, $\beta=1.2$, $R_{\rm csm}=5.5\times 10^{14}$cm, and $E_{\rm sn}=2.1\times 10^{51}$erg (red curves), a sample
not included in our training or simulated data.
@@ 2.3.1 Luminosity and photosphere property surrogates
To evaluate the performance of our trained models, we consider a range of metrics relative to an unseen training data of $30000$ simulations. Our first consideration is accuracy, across the training data, we see mean squared errors of $0.0056$, $0.06$, and $0.01$ for the luminosity, temperature, and radius surrogates, respectively, on the normalized values. Moreover, our models are also extremely fast, with each individual prediction taking $\approx 30$ms on a standard CPU. Furthermore, due to the neural network architecture’s parallel processing capabilities, we can generate predictions for large batches simultaneously. For example, all outputs for the $30,000$ test simulations are generated within $\approx 30$ms, which can provide significant benefits for inference and population studies as we will discuss later. In Fig. 1, we plot a random set of bolometric luminosities, photosphere radius and temperature from our training data (solid curves) alongside the prediction of the surrogate model (dashed curves). The bottom panel indicates the normalised residual (which we explore in more detail below), which highlights that across the bulk of times, our normalized residuals are smaller than $5\%$, this is smaller than the numerical noise in the simulations itself, particularly at early times.
We further investigate the residuals across our test data to see if there are any patterns of failure with our surrogate model. In particular, in Fig. 2, we show $1\sigma$ distribution on the normalised residuals as a function of time from the full sample of our $30000$ testing data. First, we draw attention to the luminosity residuals, which indicates a discrepancy between our surrogate model and the testing data of as large as $15\%$ at $t\sim 1$d, which could influence our results when performing inference if unaccounted for. Our residuals also highlight a peak at $\approx 100$days for all models. This broadly coincides with the rapid drop in radius and luminosity in many models around this epoch, and the residual is a product of this sharp change and the numerical noise in multiple simulations at these epochs. However, as this is only on the $\approx 5\%$ level, we consider this accurate for our purposes of inferences, as opposed to the larger error in the bolometric luminosity surrogate at early times ($t\lesssim 1$day). We note that while Fig. 2 shows the $1\sigma$ credible interval, our median normalised residuals are significantly smaller and on the order of $10^{-3}$, highlighting strong consistency between the surrogate model predictions and the training data.